OpenStudy (anonymous):

If 2 dice are tossed, what is the probability that their sum will be odd?

OpenStudy (anonymous):

Well there are 6 sides to a dice so 12 but there are only 3 odd numbers on a dice so 3/12 or 1/4

OpenStudy (anonymous):

no wait 3/6 or 1/2

OpenStudy (anonymous):

7/17?

OpenStudy (anonymous):

that doesn't even make sense @anusha.p ..the answer is 1/2

OpenStudy (anonymous):

9/21?

OpenStudy (anonymous):

Possible 11 sums \[ \{2,3,4,5,6,7,8,9,10,11,12\} \] The 5 odds are\[\{3,5,7,9,11\} \] The answer is \[ \frac 5 {11} \]

OpenStudy (anonymous):

Oh! sorry guys that whole "Sum of the odd numbers thing" my bad D: forgive me D;

OpenStudy (cwtan):

Odd+odd=even Odd+even =odd even+even=even so Odd 1st Even 2nd P(Odd+Even)= (0.5)(0.5)=0.25 Even 1st Odd 2nd P(Even+Odd)=(0.5)(0.5)=0.25 Total probability= 0.25+0.25=0.5